Mary changed a one-dollar bill for 14 coins made up of nickels and dimes. How many coins of each kind did she receive?
Answer by tutorcecilia(2144)

.05(x)+.10(14-x)=1.00
.05x+1.4-.10x=1.00
-.05x=-.40
x=8 (nickles)
Dimes = 14-x=14-8=6
.
Check by plugging the values back into the original equation.
(.05)(8)+(.10)(14-8)=1.00
1.00=1.00